# Instruction Fine-Tuning: Does Prompt Loss Matter?

### v2

@@ Abstract
We present a study analyzing the effects of prompt loss weighting (PLW) on supervised instruction fine-tuning. We recreated Stanford’s Alpaca experiment with both LLaMA 1 and LLaMA 2 and multiple instruction datasets. We found that performance of models fine-tuned on our short-completion dataset had a statistically-significant negative quadratic relationship with PLW, but performance of models fine-tuned on medium- and long-completion data did not show any relationship with PLW. I.e., prompt loss can be safely ignored for many datasets. For short-completion data, small values ($0.01-0.1$) of PLW were optimal for multiple-choice and short-generation tasks while large values ($\approx 1.0$) of PLW were optimal for long-generation tasks. We concluded that low non-zero PLW encourages models to not diverge from pre-trained model weights during training and high PLW reduces overfitting. Finally, we present a rough guide for selecting PLW values based on the completion-prompt length ratio of fine-tuning data.
@@
Mathew Huerta-Enochian EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul mathew@eq4all.co.kr
@@ 1 Introduction
Recent research in language modeling has made huge advances in training instruction-following agents. Both supervised fine-tuning and reinforcement learning have been employed to much success. However, our understanding of optimal hyperparameters and standards of practice have been slow to catch up. This research intends to contribute to hyperparameter selection for supervised instruction fine-tuning (SIFT) via an in-depth analysis of a single training hyperparameter: prompt loss weight (PLW).
The goal of SIFT is to teach a model to perform a language task given a prompted instruction. While training, model parameters are updated by optimizing for next-token maximal likelihood classification. However, it is unclear if this optimization should be performed on the entire prompt-completion sequence or only on the target completion.
Most open sourced solutions for supervised fine-tuning (SFT) either mask the prompt loss or use the entire sequence loss. On the other hand, OpenAI previously supported a prompt_loss_weight parameter that allowed users to specify PLW during SIFT.
We make the following contributions related to this parameter:
  • We show that PLW has a statistically significant negative quadratic relationship with model performance when fine-tuning on short-completion instruction data.
  • We show that PLW and prompt masking parameters can be safely ignored when fine-tuning on medium- or long-completion data, streamlining the fine-tuning process.
  • We provide evidence that the benefits of PLW for short-completion data are due to a regularizing effect and not due to previously suspected mechanisms.
  • We present a baseline for PLW selection for short-completion fine-tuning.
We provide relevant background and hypotheses in section 2 and section 3, respectively. We resent our methodology in section 4. Descriptive and regression analysis and an exploration of causal mechanisms are presented in section 5. We extend the experiment to additional datasets and present a basic model for prompt loss weight selection in section 6. We present conclusions in section 7.
@@ 2.1 Definitions
We define instruction data as one or many instances of structured text data, each containing an instruction text, an optional context or input text, and a target output text. For the rest of the paper, we use the term prompt to refer to the concatenation of instruction and input texts and the term completion to refer to the target output text. SIFT attempts to train a model to generate an appropriate completion for a given prompt.
@@ 2.2 Relevant Research and Libraries
HuggingFace’s Transformers library Wolf et al. (2020), the de facto library for training LLMs, allows users to mask select tokens when calculating loss. In Transformers, optimizing next-token prediction on prompt tokens is therefore binary—either token loss is masked (PLW $=0$) or it is unmasked (PLW $=1$).
Until recently, OpenAI supported a prompt_loss_weight parameter in their fine-tuning API, but it was officially removed as part of the v1 fine_tune API deprecation in early January, 2024. This prompt_loss_weight parameter used a default value of $0.01$ with the following parameter explanation:
“This controls how much the model tries to learn to generate the prompt (as compared to the completion which always has a weight of 1.0), and can add a stabilizing effect to training when completions are short. If prompts are extremely long (relative to completions), it may make sense to reduce this weight so as to avoid over-prioritizing learning the prompt.”
Though we could not find a study validating OpenAI’s claim or any literature that presents an analysis of PLW, there were several studies that reported parameter values for prompt_loss_weight when fine-tuning OpenAI models.
Though they do not provide their reasoning, Kozachek (2023) reported that they fine-tuned GPT-3 with a prompt_loss_weight of $0.1$. Dodgson et al. (2023) reported using the default value of $0.01$ when fine-tuning GPT models. Wang et al. (2023b) reported that a PLW of 00 performed best for them when working on the Self-Instruct framework. Interestingly, Wutschitz et al. (2023) reported hyperparameter search results for next-sentence-prediction on Elsevier data using PLWs of $0.1$ and $0.5$ and found $0.5$ to give the best results. Similar to OpenAI’s deprecated API, BLoomAI’s API supports a prompt_loss_weight parameter with a default value of 0.01.
@@
AlpacaDataShort 16.93 (13.10) 162.34 (151.69) 14.62 (10.99) 10,035,667 0.082
@@ 3 Hypotheses
Based on OpenAI’s documentation for prompt_loss_weight, we expect PLW to have a stabilizing effect on LLM SIFT when using short-completion data. Our expectation was that PLW preserves the hidden representation learned for next-token-classification when most embeddings are being used to predict tokens at the end of the sequence (as is the case when fine-tuning on short-completion data). This would represent a positive relationship between PLW and downstream performance.
However, unless the model is expected to generate prompt-like data in downstream tasks, the training and downstream sequence contexts will not match. In other words, training a model to maximize next-token-prediction on prompt tokens should be most useful for generating instruction data. Due to this train-test mismatch, we expect PLW to also have a negative influence on downstream performance.
Based on the assumption that some amount of PLW is beneficial for short-completion SIFT, we expect these two competing factors to result in a downward curved relationship between downstream performance and PLW. Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1. For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect. If $\lambda=0$, then PLW’s contribution to model performance is strictly negative. Conversely, if $\lambda=1$, then PLW contributes strictly positively to model performance.
We postulate that both $\lambda$ and the magnitude of the effect are inversely proportional to $R_{g}$ since there will be $R_{g}$ as many completion token loss terms than prompt loss terms. Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm. Rather, it would depend on numerous factors and change for each task.
We tested the relationship between PLW and model performance using three fine-tuning datasets spanning a range of $R_{g}$ values: 0.327, 7.83, and 0.082. We present the following hypotheses to be tested against each dataset.
@@
We fixed the threshold $\alpha=0.05$ significance level, and we expected to be able to reject $\mathbf{H_{0}}$ for models trained on short-completion data.
@@ 4 Methodology
We provide additional details for reproducibility in appendix C and will release our trained models on the HuggingFace Hub.
@@ 4.3 Fine-Tuning Dataset
We ran all experiments with three datasets: AlpacaData (the instruction dataset from the original Alpaca experiment), AlpacaDataCleaned, and AlpacaDataShort.
AlpacaDataCleaned Ruebsamen (2023) is a cleaned and curated version of AlpacaData that has recently been combined with data from the GPT4 LLM dataset Peng et al. (2023).
Cleaning is noted as ongoing and includes fixes for the following issues in AlpacaData: hallucinations, merged instructions, empty outputs, empty code examples, instructions to generate images, N/A outputs, inconsistent input fields, wrong answers, nonsensical instructions, and extraneous control characters.
We generated AlpacaDataShort from AlpacaDataCleaned by swapping instructions and completions for all long-completion instances and rephrasing the task as an instruction-prediction task.
For instances with empty input, we used
Predict the prompt that generated the following AI output.
as the new instruction field. For instances with non-empty input, we used
Predict the prompt that generated the below AI output given the following context. Context: <original_input>
as the new instruction field. In all cases we then used the original completion as the new input field and the original instruction as the new completion field.
Descriptive statistics for tokenized instruction, input, and completion sequences are presented in table 1. Note that AlpacaDataCleaned is strongly long-completion with an $R_{g}$ of $7.83$ while AlpacaDataShort is short-completion with an $R_{g}$ of $0.082$.
@@ 4.4 Performance Evaluation
Table 2: Benchmark tasks used for evaluation. All benchmarks were scored using Eleuther AI’s Language Model Evaluation Harness, the AlpacaEval 1 framework, and the PandaLM framework. validation splits were used when test splits were unavailable, and validation and test splits were combined for noisy benchmarks. Benchmark completion type is noted here as “MC” for multiple choice, “GS” for short generation, and “GL” for long-generation.
@@
We evaluated each model on thirteen instruction benchmarks covering both multiple choice and text generation tasks. We attempted to select benchmarks that were relatively cheap to compute and covered a range of tasks. We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH) Gao et al. (2023), Alpaca Eval 1 Li et al. (2023), and PandaLM Wang et al. (2023a, 2024). See table 2 for details on benchmark tasks.
Eleven benchmarks were run using EEH. Four benchmarks weremultiple choice tasks: ARC Challenge Clark et al. (2018), PIQA Bisk et al. (2020), TruthfulQA-MC2 Lin et al. (2022), and WinoGrande Sakaguchi et al. (2019). Seven benchmarks were short generation tasks: TruthfulQA-Gen Lin et al. (2022) and six WMT14 and WMT16 translation benchmarks Bojar et al. (2014, 2016), limited to four languages the PTLMs learned during pretraining (English, French, German, and Romanian). For WMT benchmarks, we used the instruction “Translate the following from <src_lang> to <tgt_lang>” and evaluated over both the validation and test sets to reduce variance.
Long-generation performance was evaluated using Alpaca Eval 1 and PandaLM, which are both LLM-as-a-judge frameworks.
The default auto-evaluator for Alpaca Eval 1 is GPT4, but evaluation of all training runs would be beyond the scope of this research, so we used Mixtral 8X7B Jiang et al. (2024) as an auto-evaluator. Mixtral performed the best of all available LLMs that we tested on Alpaca Eval’s evaluator test dataset Dubois et al. (2023), scoring 64.9 for agreement with human evaluators, just below Claude’s Anthropic (2023) 65.3 agreement. For reference, the default GPT4 auto-evaluator scored a 70.99 human agreement.
@@ 5.1 Performance Trends
A visualization of the simple and relative performance aggregates by wp and dataset is presented in figure 1. The simple aggregate is the unweighted mean of all task scores for given wp and dataset values. For the relative aggregate, we first min-max normalized benchmark scores over each dataset-benchmark group and then performed aggregation. The simple aggregate is dominated by large performance changes on the AlpacaEval 1 and PandaLM benchmarks. Therefore, all analysis was performed using the relative aggregate and the simple aggregate was visualized only for completeness.
[…]
Identification of trends on the six translation benchmarks was less clear due to the high level of noise relative to performance range. Aggregating scores from benchmarks translating from English to the paired language benchmarks is suggestive of an increasing relationship, but is incredibly noisy.
The performance difference across wp values for the two long-generation benchmarks was around twenty percentage points, in stark contrast to the less than two percentage point change for short-generation and multiple choice benchmarks. This suggests that PLW plays an important role in the ability to generate high quality text, and the optimal PLW for short-generation and long-generation benchmarks is clearly different.
Finally, performance of LLaMA 2 models was higher than that of LLaMA 1 models and performance of AlpacaDataCleaned models were higher than that of AlpacaData models, validating the improvements of LLaMA 2 and AlpacaDataCleaned over their predecessors.
See Appendix A for detailed figures for each benchmark.
@@ 5.2 Regression
For each data group, we fit a generalized linear mixed model (GLMM) predicting min-max transformed benchmark scores. As with the min-max transformation for the relative aggregate, benchmark scores were min-max transformed over dataset-benchmark groups.
We expected a quadratic relationship between the score and wp, so we included a second order polynomial of wp as a fixed effect. Furthermore, we knew that the performance-PLW relationship varies by benchmark and that scores were min-max normalized over each benchmark, so we used a random slope with respect to benchmark. Since we did not min-max normalize over PTLM groups and since we saw consistent improvement when using LLaMA 2, we modeled a random intercept for PTLM. This resulted in the following equation that we fit with the R library glmmTMB:
@@
This means that while wee could not reject the null hypothesis for the AlpacaData and AlpacaDataCleaned scenarios, for the AlpacaDataShort scenario, there was sufficient evidence to reject the null hypothesis in favor of the alternative hypothesis H1.
@@ 5.3 Causal Mechanism & Interpretation
To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over all five-step windows. See figure 1(a) for a boxplot of mean RSD for each model. For all dataset and PTLM factors, increasing PLW from zero led to a sharp increase in mean RSD and then a slow decrease to a minimum mean RSD at PLW $=1$. There is no obvious explanation for why training loss RSD would initially increase before decreasing.
If training loss stability was the primary factor in improved performance, we would expect stability to be best for PLW between $0.01$ and $0.1$ and for performance at PLW $=0$ to be similar with performance at PLW $=0.01$ since mean RSD at these values are similar. However, mean RSD drops by a factor of two across the PLW $\in[0.01,0.1]$ range, and performance at PLW $=0.01$ is significantly higher than the masked prompt loss scenario. Training loss mean RSD is lowest at PLW $=1$, but performance on ARC Challenge, PIQA, WinoGrande, and TruthfulQA-Gen show clear decreasing trends at this value. Furthermore, the three tasks showing positive trends at PLW $=1$ cannot be adequately explained by this factor since performance increases regardless of loss stability.
@@
We then checked if PLW was providing regularization to the weight update step, possibly improving performance by keeping weights close to the PTLM. See figure 1(b) for a visualization of weight distance from PTLM. Interestingly, for AlpacaDataShort, fine-tuned weights were closer to those of the PTLM for small values of PLW but were much farther for PLW $<0.0005$ and PLW $>0.1$. We would expect weights to change more when loss is erratic, but the range of PLW values that better preserved PTLM weights was similar to the range of increased training loss RSD. Unfortunately, we do not have an explanation for this apparent contradiction and simply conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
@@
We next explored how PLW affected training data memorization. We sampled 10,000 unique prompts from the AlpacaDataShort training set, generated completions from each prompt, and calculated corpus BLEU-4 scores. We found that for PLW from 0.0 to around 0.1, models memorized most of the training data, consistently scoring near 80 corpus BLEU. Corpus BLEU then decreased as PLW increased from 0.1. We also analyzed generation length on the AE1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW. Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
@@
The positive relationship between performance and PLW on Alpaca Eval and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
6 Predicting Prompt Loss Weight
Figure 3: Ratio-PLW Relationship
We have shown that PLW is an important hyperparameter when fine-tuning on short-completion data but is effectively irrelevant when using long-completion data. We now present a model for predicting an optimal PLW given a dataset’s $R_{g}$.
We first extended the previous experiments to two additional datasets: AlpacaDataMedium and AlpacaDataTiny to increase coverage of the parameter space. AlpacaDataMedium and AlpacaDataTiny have $R_{g}$ values of 1.0 and 0.042, respectively, and were generated similarly with AlpacaDataShort, but with different target completion-prompt ratios.
Using the fitted w-r interaction, we can now estimate an optimal PLW value for a given completion-prompt ratio $R_{g}$. See figure 3 for a visualization.
Roughly, the fitted interaction term recommends using PLW $=0.155$ for small completion-prompt length ratios ($R_{g}\leq 1$) and up to PLW $=0.278$ for a $R_{g}=1.5$ for optimal performance across all tasks. This prediction is close to our regression-predicted value of 0.242. The interaction term also confirms our conclusion that PLW is less important for data with relatively long completions.
This is a rough guide, and we recommend using it as a starting point. See appendix B for additional plots using different benchmark tasks.

### v3

@@ Abstract
We present a novel study analyzing the effects of various prompt loss token weights (PLW) for supervised instruction fine-tuning (SIFT). While prompt-masking (PLW $=0$) is common for SIFT, some fine-tuning APIs support fractional PLWs and suggest that using a small non-zero PLW can help stabilize learning when fine-tuning on short-completion data. However, there has never been a study confirming this claim, and OpenAI, a major cloud-based SIFT provider, recently removed this parameter from their fine-tuning API. We found that performance of models fine-tuned on short-completion data had a statistically-significant negative quadratic relationship with PLW. Using small values ($0.01-0.5$) of PLW produced better results on multiple-choice and short-generation benchmarks (outperforming models fine-tuned on long-completion data) while large values ($\approx 1.0$) of PLW produced better results on long-generation benchmarks. We explained this effect and verified its importance through additional experiments. This research serves as a warning to API providers about the importance of providing a PLW parameter for SIFT.
@@
Mathew Huerta-Enochian  and Seung Yong Ko EQ4ALL 11 Nonhyeon-ro 76-gil, Gangnam-gu, Seoul {mathew, stephenko}@eq4all.co.kr
@@ 1 Introduction
Recent research in language modeling has made huge advances in training instruction-following agents. Both supervised fine-tuning (SFT) and reinforcement learning (RL) have been employed to much success. However, our understanding of optimal hyperparameters and standards of practice (SOPs) have been slow to catch up. This research contributes to supervised instruction fine-tuning (SIFT) SOPs via an in-depth analysis of a single training hyperparameter: prompt loss weight (PLW).
While training, model parameters are updated by optimizing for next-token maximal likelihood classification. Most open sourced solutions for SFT either mask the prompt loss (for prefix language modeling) or use the entire sequence loss (for full language modeling) while some API providers support an explicit PLW parameter that allows users to apply fractional PLW during SIFT. The commonly-held notion is that fractional PLW helps stabilize when fine-tuning on data with short outputs. Recently, however, OpenAI quietly removed support for their prompt_loss_weight parameter.
The reason for the removal is unknown. Furthermore, to our knowledge, there has never been a proper study on the effects of PLW.
We make the following contributions:
  • We showed that PLW has a significant relationship with model performance when fine-tuning on short-completion data.
  • We showed that this relationship is due to a combination of regularizing effects and not the accepted explanation of increased training stability.
  • We provided evidence that PLW cannot be replaced by other common regularizers alone and that PLW is important for strong performance on short-generation downstream tasks.
  • We verified that PLW can be safely ignored when fine-tuning on long-completion data.
We provide relevant background and hypotheses in section 2 and 3, respectively. The main regression experiment is presented in sections 4 and 5. Supplemental experiments further validating our claims are presented in section 6. We present conclusions in section 7 followed by several appendices for additional analysis and discussion.
@@ 2.1 Definitions
We define instruction data as one or many instances of structured text data, each containing an instruction, an optional input, and a target output text. We will use the term prompt to refer to the concatenation of the instruction and input (if it exists) and the term completion to refer to the target output. The goal of SIFT is to fine-tune a model to generate an appropriate completion for a given prompt.
@@ 2.2 Relevant Research and Libraries
HuggingFace’s Transformers library Wolf et al. (2020), the de facto library for training LLMs, allows users to mask select tokens when calculating token classification loss. In Transformers, weights for next-token prediction loss is therefore binary—either token loss is masked (PLW $=0$) or it is unmasked (PLW $=1$).
As mentioned in section 1, OpenAI officially removed support for a prompt_loss_weight parameter in their fine-tuning API as part of the v1 fine_tune API deprecation in early January, 2024. This prompt_loss_weight parameter used a default value of $0.01$ with the following parameter explanation: “This controls how much the model tries to learn to generate the prompt (as compared to the completion which always has a weight of 1.0), and can add a stabilizing effect to training when completions are short. If prompts are extremely long (relative to completions), it may make sense to reduce this weight so as to avoid over-prioritizing learning the prompt.”
Though we could not find a study validating OpenAI’s claim or any literature that presents an analysis of PLW, we found several studies that reported using this parameter. Though they do not provide their reasoning, Kozachek (2023) reported that they fine-tuned GPT-3 with a prompt_loss_weight of $0.1$. Dodgson et al. (2023) reported using the default value of $0.01$ when fine-tuning GPT models. Wang et al. (2023b) reported that a PLW of 00 performed best for them when working on the Self-Instruct framework. Interestingly, Wutschitz et al. (2023) reported hyperparameter search results for next-sentence-prediction on Elsevier data using PLWs of $0.1$ and $0.5$ and found $0.5$ to give the best results. Similar to OpenAI’s deprecated API, BLoomAI’s API supports a prompt_loss_weight parameter with a default value of 0.01.
@@
AlpacaDataShort 16.93 (13.10) 162.34 (151.69) 14.62 (10.99) 10,035,667 0.08
UltraFeedback* 184.47 (268.04) 0 (-) 327.37 (291.48) 31,075,393 1.77
DatabricksDolly* 18.65 (74.55) 91.60 (250.66) 91.14 (149.15) 3,023,113 0.83
UltraFeedbackShort* 94.09 (141.94) 206.18 (211.03) 55.03 (61.00) 16,351,333 0.18
DatabricksDollyShort* 18.83 (75.29) 160.37 (270.17) 28.79 (47.49) 3,122,209 0.16
@@ 3 Hypotheses
Based on OpenAI’s explanation for prompt_loss_weight, we expected that for SIFT with short-completion data and small values of PLW, there would be a positive relationship between PLW and downstream performance. However, training a model to maximize next-token-prediction on prompt tokens should be most useful for generating instruction data, and over-prioritizing prompt token loss should have a negative influence on downstream performance.
Based on these assumptions, we would expect the two competing factors to result in a downward curved relationship between PLW and downstream performance. Limiting PLW to the range of [0, 1], we postulate that there is a critical value $\lambda$ for PLW with $10<=\lambda<=$ 1. For PLW less than $\lambda$, the positive effect dominates the negative effect and for values greater than $\lambda$, the negative effect dominates the positive effect. If $\lambda=0$, then PLW’s contribution to model performance is strictly negative, and if $\lambda=1$, then PLW contributes strictly positively to model performance. Note that $\lambda$ is not an intrinsic characteristic of the dataset, model architecture, or training algorithm. Rather, it would depend on numerous factors and change for each task.
We then made two hypotheses to test the above relationship and performed regression analysis on three fine-tuning datasets spanning a range of $R_{g}$ values: 0.08, 3.27, and 7.83.
@@
We used the standard $\alpha=0.05$ significance level, and we expected to reject $\mathbf{H_{0}}$ only for models trained on short-completion data.
@@ 4 Methodology
We provide additional details for reproducibility in appendix E and will release our trained models on HuggingFace’s Hub.
@@ 4.3 Fine-Tuning Dataset
We ran all experiments with three datasets: AlpacaData (the instruction dataset from the original Alpaca experiment), AlpacaDataCleaned Ruebsamen (2023), and AlpacaDataShort.
AlpacaDataCleaned is a cleaned and curated version of AlpacaData that has recently been combined with data from the GPT4 LLM dataset Peng et al. (2023). Cleaning is noted as ongoing and includes fixes for the following issues in AlpacaData: hallucinations, merged instructions, empty outputs, empty code examples, instructions to generate images, N/A outputs, inconsistent input fields, wrong answers, nonsensical instructions, and extraneous control characters.
We generated AlpacaDataShort from AlpacaDataCleaned by rephrasing long-completion instances as prompt-prediction task, a process we denote as prompt inversion. See Appendix A for more on prompt inversion.
Descriptive statistics for these each dataset are presented in table 1. Note that AlpacaDataCleaned is strongly long-completion with an $R_{g}$ of $7.83$ while AlpacaDataShort is short-completion with an $R_{g}$ of $0.082$.
@@ 4.4 Performance Evaluation
Table 2: Evaluation benchmarks. Validation splits were used when test splits were unavailable, and validation and test splits were combined for noisy benchmarks. Benchmark completion type is noted here as “MC” for multiple choice, “GS” for short generation, and “GL” for long-generation.
@@
We evaluated each model on thirteen instruction benchmarks covering multiple choice and text generation tasks. We selected benchmarks that were relatively cheap to compute and covered a range of tasks. We used three evaluation frameworks: EleutherAI’s Language Model Evaluation Harness (EEH) Gao et al. (2023b), AlpacaEval 1 Li et al. (2023), and PandaLM Wang et al. (2023a, 2024). See table 2 for details on benchmark tasks.
Eleven benchmarks were run using EEH. Four of these were multiple choice tasks: ARC Challenge Clark et al. (2018), PIQA Bisk et al. (2020), TruthfulQA-MC2 Lin et al. (2022), and WinoGrande Sakaguchi et al. (2019). Seven were short generation tasks: TruthfulQA-Gen Lin et al. (2022) and six WMT14 and WMT16 translation benchmarks Bojar et al. (2014, 2016), limited to four languages the PTLMs saw during pretraining (English, French, German, and Romanian). For WMT benchmarks, we used the zero-shot instruction “Translate the following from <src_lang> to <tgt_lang>” and evaluated over both the validation and test sets to reduce variance.
Long-generation performance was evaluated using AlpacaEval 1 and PandaLM, which are both LLM-as-a-judge frameworks. The default auto-evaluator for AlpacaEval 1 is GPT4, but using paid APIs would beyond the scope of this research, so we used Mixtral 8X7B Jiang et al. (2024) as an auto-evaluator. Mixtral performed the best of all open-source LLMs that we tested on AlpacaEval’s evaluator test dataset Dubois et al. (2023), with 64.9% agreement with human evaluators. For reference, Claude Anthropic (2023) has 65.3% and GPT4 has 70.99% human agreement.
@@ 5 Results and Discussion
A visualization of the simple and min-max scaled relative performance aggregates by wp and dataset is presented in figure 1. Note that all analysis was performed using the relative aggregate. The simple aggregate is dominated by large performance changes on a few benchmarks and is included only for completeness.
@@ 5.1 Performance Trends
There are several qualitative performance trends that are of interest. For more thorough discussion and task-specific benchmark plots, see Appendix B.
A group of four benchmarks (Arc Challenge, PIQA, TruthfulQA-Gen, and WinoGrande) clearly show the expected negative quadratic relationship between wp and performance of AlpacaDataShort models. Notably, AlpacaDataShort models outperform AlpacaData and AlpacaDataCleaned on this group given optimal PLW tuning. On the long-generation benchmarks, however, AlpacaDataShort models show a steadily-increasing trend. For these seven benchmarks and AlpacaDataShort models, using prompt loss masking (wp$=0$) almost always produced the worst scores.
The wp-based performance increase was around twenty percentage points for both long-generation benchmarks. Contrasted with the less than two percentage point change for short-generation and multiple choice benchmarks, and it is evident that the optimal PLW for short-generation and long-generation benchmarks is different.
Clear qualitative performance trends for the AlpacaData and AlpacaDataCleaned and for any model on the six translation benchmarks could not be identified.
@@ 5.2 Regression
For each data group, we fit a generalized linear mixed model (GLMM) with the relative aggregate benchmark scores as the response variable.
We expected a quadratic relationship between the score and wp, so we included a second order polynomial of wp as a fixed effect. Furthermore, we knew that the PLW-performance relationship varies by benchmark and since scores were min-max normalized over each benchmark, we used a random slope (and no intercept) with respect to benchmark. Since we did not min-max normalize over PTLM groups and since we saw consistent improvement when using LLaMA 2, we modeled a random intercept for PTLM. This resulted in the following equation that we fit with the R library glmmTMB:
@@
This means that while we could not reject the null hypothesis for the AlpacaData and AlpacaDataCleaned scenarios, for the AlpacaDataShort scenario, there was sufficient evidence to reject the null hypothesis in favor of the alternative hypothesis H1.
@@ 5.3 Causal Mechanism & Interpretation
To identify possible causal mechanisms, we first investigated the effects of PLW on training stability by analyzing training loss relative standard deviation (RSD) over five-step windows. See figure 2(a) for a boxplot of mean RSD for each model. For all dataset and PTLM factors, increasing PLW from zero led to a sharp increase in mean RSD and then a slow decrease to a minimum mean RSD at PLW $=1$. There is no obvious explanation for why training loss RSD would increase for small PLW before decreasing for large PLW.
If training loss stability was the primary factor in improved performance, we would expect RSD to be lowest for PLW between $0.01$ and $0.5$ (or even between $0.01$ and $0.1$ based on short-generation benchmarks) and for performance at PLW $=0$ to be similar with performance at PLW $=0.01$ since mean RSD at these values are similar. However, mean RSD drops by a factor of two across the PLW $\in[0.01,0.1]$ range, and performance at PLW $=0.01$ is significantly higher than the masked prompt loss scenario. Training loss mean RSD is lowest at PLW $=1$, but performance on ARC Challenge, PIQA, WinoGrande, and TruthfulQA-Gen show clear decreasing trends at this value. Furthermore, the three tasks showing positive trends at PLW $=1$ cannot be adequately explained by this factor since performance increases regardless of loss stability.
@@
We then checked if PLW was providing regularization to the weight update step, possibly improving performance by keeping weights close to the PTLM. See figure 2(b) for a visualization of weight distance from PTLM. Interestingly, for AlpacaDataShort, fine-tuned weights were closer to those of the PTLM for small values of PLW but were much farther for PLW $<0.0005$ and PLW $>0.1$. We would expect weights to change more when loss is erratic, but the range of PLW values that better preserved PTLM weights was similar to the range of increased training loss RSD. This is an interesting result, and we conclude that PTLM model weights were better preserved for small non-zero PLW despite high loss instability.
We next explored how PLW affected training data memorization. We sampled 10,000 unique prompts from the AlpacaDataShort training set, generated completions from each prompt, and calculated corpus BLEU-4 scores.
@@
We found that for PLW from 0.0 to around 0.1, models memorized most of the training data, consistently scoring near 80 corpus BLEU. Corpus BLEU then decreased as PLW increased from 0.1. We also analyzed generation length on the AlpacaEval 1 test set which showed high deviations in generation length for the high loss RSD PLW values for LLaMA 2 only and a generally increasing length trend with PLW. Considering that AlpacaDataShort contains short completions, this strongly suggests that PLW contributes a regularizing effect and helps decrease overfitting.
@@
The positive relationship between performance and PLW on AlpacaEval 1 and PandaLM as opposed to the negative quadratic relationship seen on several of the other benchmarks can then be explained by a tradeoff between longer generations and decreased overfitting for high PLW and better preservation of PTLM weights for low PLW.
6 Supplemental Experiments
In this section, we present supplemental PLW experiments that suggest that PLW cannot be replaced by alternative regularization techniques for SIFT and that the effects of PLW extend to other short-completion datasets. Since the translation benchmarks showed high levels of noise and unclear correlation with PLW in the main experiment, they were not included in supplemental experiments.
6.1 Alternative Regularizers
LLaMA Regularization Aggregate
Simple Relative
1 Weight Decay 0.507 0.795
Minkowski Metric 0.505 0.812
Dropout 0.500 0.783
Label Smoothing 0.500 0.785
PLW (Ours) 0.506 0.855
2 Weight Decay 0.537 0.773
Minkowski Metric 0.538 0.772
Dropout 0.541 0.805
Label Smoothing 0.538 0.837
PLW (Ours) 0.537 0.894
Table 4: Supplemental comparison of AlpacaDataShort models fine-tuned with various regularizers. High scores are in bold. As explained in section 5, the relative aggregate should be used for analysis, and the simple aggregate is provided for reference only.
To investigate if the effects of PLW on short-completion SIFT can be emulated with other common regularization techniques, we repeated the AlpacaDataShort training runs for PLW $=0$ and PLW $=1$ while applying various regularizers. We tested weight decay, minimizing the Minkowski distance between PTLM weights and learned weights, dropout, and label smoothing. See Appendix C.1 for additional visualizations.
Note that we do not compare PLW with KL-divergence. While using KL-divergence as a regularizing loss is common for SFT in general and is used in RL-based LLM alignment (Stiennon et al., 2020; Korbak et al., 2022; Gao et al., 2023a), collecting embeddings for any non-trivial dataset to be used for LLM SIFT presents a huge computation and memory overhead.
As can be seen in the results in table 4, relative aggregate scores were higher for models fine-tuned with fractional PLWs. This suggests that PLW provides a unique regularizing effect that cannot be easily replaced with other regularizers.
6.2 Alternative Datasets
(a) Original Datasets
(b) Prompt-Inverted Datasets
Figure 3: Relative aggregate scores showing the effects of PLW for SIFT on alternative datasets. (a) UltraFeedbackCleaned and DatabricksDolly models. (b) UltraFeedbackShort and DatabricksDollyShort models.
We repeated our main training experiment using two additional datasets, a cleaned and binarized version of UltraFeedback (Cui et al. (2023); denoted UltraFeedbackCleaned) and databricks-dolly-15k (Conover et al., 2023; denoted DatabricksDolly). We also trained on prompt-inverted versions of both datasets, denoted UltraFeedbackShort and DatabricksDollyShort, respectively, and expanded analysis to include LLaMA 3 8B AI@Meta (2024) as a PTLM. See length statistics for these four datasets in table 1 and appendix C.2 for additional visualizations.
The combined relative aggregate scores for UltraFeedbackCleaned and DatabricksDolly and for UltraFeedbackShort and DatabricksDollyShort are visualized in figure 3. Visual inspection suggests that PLW affected learning for both groups of datasets. For the shortened variants, performance appears to have the same negative quadratic relationship with PLW as AlpacaDataShort models do in the main experiment. The relationship between PLW and performance for models fine-tuned on the original datasets is less clear but appears to have a generally increasing trend, unlike the long-completion datasets in the main experiment. We believe this is due to both UltraFeedbackCleaned and DatabricksDolly having much smaller completion-prompt length ratios $R_{g}$ than the AlpacaData and AlpacaDataCleaned datasets.